OpenStudy (anonymous):

what is sqrt{2} ? and 8/9 ? *natural # *whole # *integer *rational # *irrational # *real #

OpenStudy (anonymous):

the square-root of 2 is irrational

OpenStudy (anonymous):

And I think 8/9 is a rational number. Also this is a pretty good math website, http://www.purplemath.com/modules/numtypes.htm

OpenStudy (accessdenied):

Some definitions: \[ \quad \textbf{Natural: } \text{the "counting" numbers} \\ \qquad \{1, 2, 3, ...\} \\ \quad \textbf{Whole: } \text{Natural numbers including 0} \\ \qquad \{0, 1, 2, 3, ...\} \\ \quad \textbf{Integers: } \text{The whole numbers and their opposites} \\ \qquad \{..., -2, -1, 0, 1, ...\} \\ \quad \textbf{Rationals: } \text{The set of all fractions of two integers} \\ \qquad \{\frac{p}{q}, p \text{ and } q \text{ are integers} \} \\ \quad \textbf{Irrationals: } \text{Basically not rational / cannot be expressed as fraction .} \\ \quad \textbf{Reals: } \text{The entire set of rationals and irrationals} \\ \] Yeah, \(\sqrt{2}\) is irrational and real. \( \frac{8}{9} \) is rational and real.

OpenStudy (anonymous):

"Whole" is not an accepted mathematical definition....